The average molar mass of a mixture of methane and ethane is equal to 30 g / mol. The mass of methane in the mixture is 64 g. Determine the amount of ethane in this mixture.
1
Expert's answer
2020-03-23T10:15:57-0400

M(C2H6) = 30.07 g / mol; M(CH4) = 16.04 g / mol;

30.07x + 16.04(1 - x) = 30

x = 0.995

m(C2H6) = 30.07 × ((64 / 16.04) / 0.005) = 23996 g.

Answer: 23996 g.